3. Find five rational numbers between 35 and 45 .

Grade:12th pass

1 Answers

Pawan Prajapati
3 years ago
Here may be infinite rational number between any two rational number. But here we have to determine only five number between 35 and 45 . To get five rational number between 35 and 45 in easiest way we write 35 and 45 as follows 3 = 3#(5+1) = 3#6 = 18 5 5 #(5+1) 5 # 6 30 and 4 = 4 #(5+1) = 4 # 6 = 24 5 5 #(5+1) 5 # 6 30 Five rational numbers between 35 and 45 are 1930 , 2030, 3021 , 3022, 2330 . There can be other set of rational numbers also.